### v1

# Robustness and Cybersecurity in the EU Artificial Intelligence Act

Abstract.
The EU Artificial Intelligence Act (AIA) establishes different legal principles for different types of AI systems. While prior work has sought to clarify some of these principles, little attention has been paid to robustness and cybersecurity. This paper aims to fill this gap. We identify legal challenges and shortcomings in provisions related to robustness and cybersecurity for high-risk AI systems (Art. 15 AIA) and general-purpose AI models (Art. 55 AIA). We show that robustness and cybersecurity demand resilience against performance disruptions. Furthermore, we assess potential challenges in implementing these provisions in light of recent advancements in the machine learning (ML) literature. Our analysis informs efforts to develop harmonized standards, guidelines by the European Commission, as well as benchmarks and measurement methodologies under Art. 15(2) AIA. With this, we seek to bridge the gap between legal terminology and ML research, fostering a better alignment between research and implementation efforts.
Robustness, Cybersecurity, AIA, Regulation, European Union, EU law
copyright: nonejournalyear: 2025
1. Introduction
The European Union (EU) recently adopted the Artificial Intelligence Act (AIA) [^1] EU Regulation 2024/1689, 12.7.2025. which creates a legal framework for the development, deployment, and use of “human-centered and trustworthy artificial intelligence (AI)” (Art. 1 AIA). The AIA outlines desirable “ethical principles” of AI systems (Rec. (27)) and, i.a., imposes some of these as legally binding requirements for high-risk AI systems (HRAIS), e.g, AI systems intended to be used to take university admission decisions, or to evaluate individuals’ creditworthiness, and for general-purpose AI models (GPAIMs), e.g., multimodal large language models. While the AIA is recognized as being one of the first legally binding regulatory frameworks for AI (Chee and Hummel, 2024), it has faced criticism for its imprecise and incoherent terminology (Laux et al., 2024; Bomhard and Siglmüller, 2024), which will complicate its practical implementation. Previous work has examined the AIA and its legislative history to clarify terms like explainability (Bordt et al., 2022; Vitali, 2022; Pavlidis, 2024) and fairness (Deck et al., 2024). So far, little attention has been paid to robustness and cybersecurity. Only AI systems classified as high-risk (HRAIS) must meet the robustness and cybersecurity requirements set out in Art. 15 AIA. This paper thus focuses on requirements for HRAIS. To provide a clearer understanding, we compare these requirements with requirements for specific AI models, namely for GPAIMs with systemic risk, in Art. 55 AIA.
Technical solutions to ensure the robustness and cybersecurity of AI systems are often developed within the ML domain. Therefore, it is essential to inform ML research about the legal requirements to ensure compliance with the AIA. However, the vagueness of requirements for cybersecurity and robustness under the AIA makes it challenging to inform ML practitioners about the specific legal requirements to further the development of solutions that can ensure compliance with the AIA. A common understanding between technical and legal domains can be facilitated through technical standards. While the AIA sets out general rules, technical standards specify these rules in detail. Standards are technical specifications designed to provide voluntary technical or quality specifications for current or future products, processes or services. [^2] Art. 1, 2(1) EU Regulation 1025/2012, OJ L 316, 14.11.2012. They prescribe technical requirements, including characteristics such as quality or performance levels, terminology, and test methods. [^3] Art. 2(4)(a) and (c) ibid. Standards have long been integral to EU product legislation under the New Legislative Framework, upon which the AIA is built (Gorywoda, 2009). If approved by the EU Commission, technical standards become harmonized technical standards, which grants a presumption of conformity to products or processes that adhere to them. Consequently, compliance with these standards is deemed to fulfill the requirements of the AIA, thereby incentivizing providers to adopt them (Art. 40 AIA). The development of harmonized technical standards for the AIA has been initiated by the EU Commission and is expected to be completed in the next few years. In addition, the EU Commission is also tasked with developing additional guidelines on the practical implementation on the application of Art. 15 AIA (Art. 96(1)(a) AIA).
In this paper, we make the following contributions:
  • We analyze and explain the legal requirements related to robustness and cybersecurity in the AIA, identify related shortcomings, and offer possible solutions for some of these shortcomings.
  • We evaluate these findings in relation to their practical implementability. This aims to inform the standardization process, the development of guidelines by the EU Commission, as well as the benchmark and measurement methodologies referred to in Art. 15(2) AIA.
  • We connect the legal requirements for robustness and cybersecurity to ML terminology, aiming to inform ML research and ensure that technical solutions are conducive to legal compliance.
This paper is structured as follows: Section 2 provides a short background on robustness and cybersecurity in the ML literature. Section 3 provides an introduction to the AIA and Art. 15 AIA. Section 5 analyzes the requirements outlined in Art. 15 AIA for HRAIS, addressing both general challenges pertinent to robustness and cybersecurity, as well as specific issues related to each requirement. Section 6 examines the requirements in Art. 55 AIA relevant to GPAIMs with systemic risk. Section 7 concludes with a summary and recommendations for future research.
2. An ML Perspective on Robustness and Cybersecurity
ML research on robustness focuses on mitigating undesired changes in model outputs when deploying models in real world scenarios (Schwinn et al., 2022). This issue is explored across different applications such as computer vision (Drenkow et al., 2021; Taori et al., 2020; Dong et al., 2020) and natural language processing (La Malfa and Kwiatkowska, 2022; Chang et al., 2021). Unintended changes in model outputs can occur due to adversarial or non-adversarial factors affecting the ML model, its input (test) data, or its training data (Cheng et al., 2024; Tocchetti et al., 2024).
Perturbations of input (test) data often present a difficult challenge (see Figure 2). While a model’s output may be as expected when using “safe” test data from the original population distribution, unintended changes can occur when perturbed examples are provided as input to the ML model.
Non-adversarial (or natural) robustness often addresses changes in ML model outputs due to distribution shifts in input data (Tocchetti et al., 2024; Gojić et al., 2023). These changes occur when the distribution from which the test data is sampled differs from that of the training data (Drenkow et al., 2021; Taori et al., 2020). For instance, alterations in data collection methods, such as upgrading to a new X-ray machine, can modify the format or presentation of images (Glocker et al., 2019; Castro et al., 2020). Importantly, distribution shifts can also result from feedback loops, where the ML model’s outputs influence the data distribution, creating a cycle from the model’s output back to its input (D’Amour et al., 2020; Zhang et al., 2020). Such an effect can be found, for example, in movie recommendation systems, where user’s preferences change over time in response to the ML system’s suggestions, thereby influencing future recommendations (Perdomo et al., 2020). Other forms of research on non-adversarial robustness investigate the robustness of ML models to noise, which frequently occurs in real-world data sets (Sáez et al., 2016; Olmin and Lindsten, 2022).
Adversarial robustness refers to the study and mitigation of model evasion attacks using adversarial examples. These are data samples typically drawn from the original population distribution and then modified by an adversary—often in ways that are difficult or even impossible to detect through human oversight—with the intent of altering a model’s output (Szegedy et al., 2013). This phenomenon can occur across various models and data types. For instance, in the image domain, small pixel perturbations in input images can lead to significant changes in a model’s output (Szegedy et al., 2013). In a broader sense, adversarial robustness also encompasses the study and mitigation of other forms of adversarial attacks that attempt to extract the model or reconstruct or perturb the training data set (Nicolae et al., 2018; Chen et al., 2017a).
From a technical standpoint, adversarial robustness is one aspect of cybersecurity. Research on cybersecurity focuses on developing defenses that protect computer systems from attacks compromising their confidentiality, integrity, or availability (Dasgupta et al., 2022). This encompasses aspects like data storage, information access and modification, and secure data transmission over networks (Sarker et al., 2021). Unlike robustness, cybersecurity is not a stand-alone concept in ML, but is discussed more broadly as both a tool for ensuring cybersecurity and a potential source of cybersecurity risks. [^4] E.g., NeurIPS’18 Workshop on Security in ML (Papernot et al., 2018), ICML’22 Workshop on ML for Cybersecurity (Applebaum et al., 2022). ML algorithms can be employed to detect and mitigate cybersecurity threats (Sarker et al., 2021), but can also introduce specific vulnerabilities that adversaries may exploit, such as data poisoning or adversarial attacks (Roshanaei et al., 2024; Rosenberg et al., 2021).
Figure 1. Robustness problems. Outputs may be as expected (✓) with “safe” test data from the original distribution; unintended changes (✗) can occur with adversarial or non-adversarial (shifted) inputs.
Figure 2. Technical solutions to cybersecurity can be found, i.a., in ML research on adversarial robustness; technical solutions to robustness can be found, i.a., in the ML research on non-adversarial robustness.
[…]
3. Background on the AIA and Art. 15 AIA
AIA
The AIA creates harmonized rules for certain AI systems in order to incentivize the use of such systems in the EU market and prevent regulatory fragmentation between member states (Rec. (1)). It is formally structured into recitals (Rec.), articles (Art.), and annexes. Recitals are not strictly legally binding and outline the rationale behind the articles, articles delineate specific binding obligations, and the annexes provide additional details and specifications to support the articles  (Klimas and Vaiciukaite, 2008). [^5] From now on, whenever we cite recitals, we refer to those in the AIA without explicitly indicating it. Art. 3(1) AIA defines an AI system as “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs […] that can influence physical or virtual environments”. These AI systems are regulated differently based on their perceived risk level (Bomhard and Siglmüller, 2024; Sioli, 2021): Those posing unacceptable risks, such as social scoring, are prohibited or subject to qualified prohibitions; HRAIS, such as those used in medical devices, are allowed but must comply with certain requirements and undergo pre-assessment; other AI systems are subject only to specific transparency and information obligations. Among these categories, only HRAIS must fulfill the robustness and cybersecurity requirements under Art. 15 AIA. According to Art. 16(a) AIA, providers of HRAIS must ensure compliance with these requirements. An AI system is considered a HRAIS if it is either a safety component of a product or a product itself regulated under specific legislation, such as medical devices, machinery, or toys (Art. 6(1) AIA, Annex I), or if it poses a significant risk of harm to the health, safety, or fundamental rights of individuals in specific areas, such as education, employment, or law enforcement (Art. 6(2) and (3) AIA, Annex III).
In addition to AI systems, the AIA establishes a separate regime of legal requirements in chapter V of the AIA for a very specific type of AI models, namely GPAIM (e.g., multimodal large language models, see also Section 6). GPAI models are AI models that can perform tasks that they were not originally trained for (Gutierrez et al., 2023), such as large language models (OpenAI, 2023; Gemini Team et al., 2023), or large text-to-image models (Ramesh et al., 2022). Other types of AI models are not regulated and mentioned in the AIA. These are models that are created with a specific objective and can only accomplish tasks they are trained to perform (e.g., translation, classification).
The AIA does not directly define specific technical requirements. Instead, it sets out ’essential requirements’ that AI systems must comply with, and which are concretized by so-called technical standards. The regulatory concept of relying on standardization is a well-established process in EU product legislation and is referred to as the New Legislative Framework (Gorywoda, 2009). It traditionally involves the participation of stakeholders, such as providers of AI systems. The European Commission has already issued a standardization request to the European Committee for Standardisation and the European Committee for Electrotechnical Standardisation to develop standards for the AIA until 30 April 2025.666C(2023)3215 - Standardisation request M/593. Specifically, Annex I enlists ‘robustness’ and ‘cybersecurity’ specifications for AI systems as standardisation deliverables to be developed. Once accepted by the European Commission, these technical standards become ’harmonized’ standards and compliance with the AIA will be presumed if providers adhere to them (Art. 40 AIA).
Art. 15 AIA
Art. 15(1) AIA requires that HRAIS “shall be designed and developed in such a way that they achieve an appropriate level of accuracy, robustness, and cybersecurity, and that they perform consistently in those respects throughout their lifecycle”. The provision outlines specific product-related requirements for AI systems to ensure they are trustworthy. Art. 15(4) AIA mandates that HRAIS exhibit resilience “regarding errors, faults or inconsistencies that may occur within the system or the environment in which the system operates, in particular due to their interaction with natural persons or other systems”. Additionally, Art. 15(5) AIA requires HRAIS to be “resilient against attempts by unauthorised third parties to alter their use, outputs or performance by exploiting system vulnerabilities”. Art. 15(2) AIA requires the EU Commission, together with other relevant stakeholders, to encourage the development of benchmarks and measurement methods for assessing accuracy, robustness, and other performance metrics.
4. The Purpose of Art. 15 AIA
Art. 15(1) AIA states that high-risk AI systems must achieve an “appropriate level of accuracy, robustness and cybersecurity” and must function consistently in this respect throughout their lifecycle. It is one of seven provisions in the AIA that sets specific requirements for high-risk AI systems. These requirements aim to ensure product safety and a “consistent and high level of protection of public interests as regards health, safety and fundamental rights” (Rec. (7)). Particularly, Art. 15(1) AIA outlines specific product-related requirements for AI systems, detailing how they must be “designed and developed” to achieve trustworthy AI. This is emphasized by the 2019 Ethics Guidelines for Trustworthy AI developed by the AI Independent High-level Expert Group (AI IHEG) on AI (AI IHEG, 2019) appointed by the European Commission which can be seen as the conceptual basis of the AIA (Rec. (27)). These guidelines state that technical robustness ensures that AI systems “reliably behave as intended while minimizing unintentional and unexpected harm, and preventing unacceptable harm”.
The question arises as to why it is even necessary to regulate accuracy, robustness and cybersecurity of AI systems. One could assume that it is in the best interest of an economic actor to fulfill these requirements in the best possible way to gain a market advantage. The European Commission’s impact assessment of the AI Act recognizes this thought with respect to accuracy and robustness, stating that “an economic operator […] would anyway have to ensure that their product actually works” (European Commission, 2021, Annex 4). However, the impact assessment clarifies that “is important that these requirements are included in the regulatory framework so that substandard operators need to improve their procedures” (European Commission, 2021, Annex 4).
Therefore, the purpose of the specific requirements set out in Art. 15 AIA is to achieve the overarching objective of ensuring the trustworthiness of AI systems (see Art. 1(1) AIA) and to advance the cybersecurity agenda of the EU. [^7] COM/2010/245 final/2, COM/2021/118 final.
5. Requirements for High-Risk AI Systems
In this section, we provide an analysis of the overarching challenges of implementing Art. 15 AIA (Section 5.1), followed by a discussion regarding the robustness requirement in Art. 15(4) AIA (Section 5.2) and the cybersecurity requirement in Art. 15(5) AIA (Section 5.3).
5.1. General Challenges of Art. 15 AIA
We identify four legal challenges related to Art. 15 AIA that may arise in its practical implementation. First, there is no clear delineation of the legal terms of robustness and cybersecurity and its counterparts in ML literature. Second, while the AIA mandates compliance for entire AI systems, the ML literature primarily focuses on models, which may pose practical challenges for implementation. Third, while accuracy is specified as a requirement in Art. 15 AIA, the provision does not clarify its role in measuring robustness and cybersecurity. Fourth, the terms ’lifecycle’ and ’consistent’ performance are not defined, leaving ambiguity about how such performance can be practically ensured.
Robustness and Cybersecurity.
The robustness requirement in Art. 15(4) AIA addresses “errors, faults, or inconsistencies” that may inadvertently occur as the system interacts with its real-world environment. In contrast, the cybersecurity requirement in Art. 15(5) AIA targets deliberate attempts “to alter the use, outputs, or performance” of an AI system “by malicious third parties exploiting the system’s vulnerabilities”. Both robustness and cybersecurity requirements aim to ensure that HRAIS perform consistently and are resilient against any factors that might compromise this performance. They, however, address different threats to consistent performance: robustness requires protection against unintentional causes, whereas cybersecurity protects against intentional actions. While robustness is a new term in EU legislation and not explicitly defined in the AIA, the term cybersecurity has already been defined in the EU Cybersecurity Act (CSA) [^8] Regulation (EU) 2019/881, OJ L 151, 7.6.2019.. Art. 2(1) CSA provides a broad definition of cybersecurity, covering all “the activities necessary to protect network and information systems, the users of such systems, and other persons affected by cyberthreats”. According to the CSA, a cyber threat is any potential circumstance, event or action that could damage, disrupt or otherwise adversely impact network and information systems, the users of such systems and other persons (Art. 2(8) CSA). Importantly, the CSA does not distinguish between intentional or unintentional cyberthreats as causes of harm. Rather, both are explicitly included in the scope of the CSA (see, e.g., Art. 51(1)(a) and (b) CSA). However, the AIA artificially splits the CSA’s concept of cybersecurity by designating unintentional causes as a matter of robustness and restricting cybersecurity to intentional actions. This creates a conflict when aligning the AIA’s requirements with the CSA’s definition of cybersecurity that may lead to regulatory ambiguity. Specifically, Art. 42(2) AIA considers HRAIS with CSA certification or conformity declarations as compliant with cybersecurity requirements in Art. 15 AIA. [^9] Note that this holds only true in so far as the cybersecurity certificate or statement of conformity or parts thereof cover those requirements in Art. 15 AIA. This suggests that the CSA definition of cybersecurity applies to the AIA, even though it inherently covers both types of causes.
We explore how these legal terms could be understood within the ML domain proposing a simple model as an explanatory heuristic (see Figure 2). In ML, robustness refers to maintaining consistent model performance in real-world scenarios (Schwinn et al., 2022). ML research distinguishes between different types of robustness. Non-adversarial robustness in ML refers to a model’s ability to maintain performance despite data shifts or noise (Tocchetti et al., 2024; Gojić et al., 2023; Sáez et al., 2016; Olmin and Lindsten, 2022). This aligns with the legal term robustness in the AIA. Adversarial robustness in ML refers to the model’s resistance to intentional perturbations aimed at altering predictions (Szegedy et al., 2013). This aspect aligns more with the legal concept of cybersecurity. The cybersecurity requirement in Art. 15(5) AIA aims to ensure AI systems’ integrity, confidentiality, and availability, protecting them from threats like unauthorized access, adversarial manipulation, data modification, Denial-of-Service attacks, and theft of sensitive information (e.g., model weights). However, other scenarios within the ML domain may also fall under the relevant legal terms. For example, language model jailbreaks exploit AI vulnerabilities to bypass safety constraints (Wei et al., 2024; Vassilev et al., 2024). This aligns more closely with the notion of cybersecurity in protecting against misuse of AI systems.
Our findings are supported by an historic analysis of the legislation process. As outlined in Section 4, the AIA builds on the Ethics Guidelines for Trustworthy AI (AI IHEG, 2019). In the guidelines, the principle of ’technical robustness and safety’ includes resilience against attacks, but does not mention cybersecurity. The White Paper on Artificial Intelligence (European Commission, 2020), which elaborates on these guidelines, still lists resilience to attacks against AI systems under “robustness and accuracy” without differentiating those terms from cybersecurity. The first official draft of the AIA by the European Commission [^10] COM/2021/206 final. was the first official document to distinguish between these three terms and assigned “resilience against attacks” to cybersecurity rather than robustness. Rec. (27), which refers to the IHGE guidelines, seems to be a remnant of this development process. It demands under the term ‘technical robustness’ that AI systems should be resilient “against attempts to alter the use or performance of the AI system”, essentially asking for adversarial robustness.
System vs. Model.
The AIA regulates AI systems, but not AI models, with the only exception being GPAIMs. ML research, in contrast, often focuses on developing technical solutions for ML models. This raises the question of whether solely relying on technical solutions for ML models is enough to ensure the compliance of a HRAIS with Art. 15 AIA—or whether additional measures are needed. Rec. (97) specifies that an AI model is an essential component of an AI system. [^11] Although Rec. (97) specifically refers to GPAIMs, the wording suggests that the statement about the relationship between AI systems and AI models is of a general nature. Additional components can include, i.a., user interfaces, sensors, databases, network communication components, or pre- and post-processing mechanisms for model in- and outputs (Rec. (97), (Junklewitz et al., 2023)). All these individual components should contribute to the overall robustness of the AI system, particularly in scenarios where some components may fail. This is illustrated by Art. 15(4)(ii) AIA, which states that robustness may be ensured through technical redundancy solutions, including “back-up or contingency plans”. Furthermore, Art. 15(5)(iii) AIA stipulates that the cybersecurity of AI systems shall be achieved through technical solutions that, “where appropriate”, target training data, pre-trained components, the AI model or its inputs. This binding provision suggests that at least these different components of the AI system are required to be assessed individually for their appropriateness in mitigating cybersecurity attacks. Thus, Art. 15 AIA should not be understood as requiring a single, unified assessment of the requirements. Instead, it must be interpreted as mandating that each component, including one or more ML models, be assessed individually. The assessment of the AI system’s overall performance is then derived from an aggregation of the individual performance results (Kumar and Anderson, 2023). This requires an interdisciplinary approach that draws on expertise from fields such as ML, engineering, and human-computer interaction. To establish a common understanding, it can prove beneficial to formally describe the evaluation process of an entire AI system, including potential challenges, such as interdependencies of technical solutions.
Role of Accuracy.
Art. 15(1) AIA mandates that HRAIS shall “achieve an appropriate level of accuracy”. This is important because trade-offs between different desiderata can exist, such as between robustness and accuracy (see Appendix E). While accuracy is not defined in the AIA, Annex IV No. 3 AIA states that accuracy is an indicator of the capabilities and performance limits of an AI system. Accordingly, accuracy should be measured in at least two ways: i) separately for “specific persons or groups of persons on which the system is intended to be used” [^12] This links to fairness ML literature on diverging error rates for different sensitive groups (Mitchell et al., 2021; Chouldechova and G’Sell, 2017)., and ii) the overall expected accuracy for the “intended purpose” of the AI system. In ML, the metric accuracy typically describes the overall proportion of correct predictions out of the total number of predictions made (Carvalho et al., 2019). However, the term can also describe the objective of “good performance” of an AI system and, depending on its specific purpose, can also be evaluated using different metrics, such as utility (Corbett-Davies et al., 2017) and f1-score (Sokolova et al., 2006). Art. 15(3) AIA explicitly references ‘accuracy and the relevant accuracy metrics’, indicating that accuracy is understood as an objective that can be measured with various metrics, leaving the choice of the relevant metric to the provider. The selection of the metric should consider various factors, including the specific purposes of the ML model, dataset-specific circumstances (e.g., imbalanced data) and the particular model type (e.g., classification, regression). Technical standards and guidelines by the EU Commission should clarify how AI systems’ accuracy should be measured.
In ML, robustness is often measured using an accuracy metric. Typically, this involves comparing the accuracy (or error rates) evaluated on an unperturbed dataset from the original distribution with the accuracy on a perturbed test set (e.g., sampled from the shifted distribution or containing adversarial samples) (Taori et al., 2020; Hendrycks et al., 2021; Goodfellow et al., 2015). The smaller the difference between these two accuracy results, the better the robustness. The choice of the accuracy metric thus has an impact on the measurement of robustness. As a result, the ML model may appear more robust under some accuracy metrics than others. The selection of favorable metrics has been studied in fair ML under the term fairness hacking (Meding and Hagendorff, 2024; Simson et al., 2024; Black et al., 2024). Without entering into the debate, we note that there is an ongoing discussion in the ML literature about the existence and characteristics of a trade-off between robustness and accuracy. While some research showed that enhancing robustness leads to a drop in test accuracy (Zhang et al., 2019; Rade and Moosavi-Dezfooli, 2022; Tsipras et al., 2019), others believe that robustness and accuracy are not conflicting goals and can be achieved simultaneously (Yang et al., 2020; Raghunathan et al., 2020). Technical standards and guidelines by the EU Commission should provide instructions on how AI system providers should choose an appropriate ‘accuracy’ measure, especially when it is used to assess robustness in subsequent steps.
Consistent Performance Throughout the Lifecycle.
AI systems must perform “consistently” in terms of accuracy, robustness, and cybersecurity “throughout their lifecycle” (Art. 15(1) AIA). Performance is the “ability of an AI system to achieve its intended purpose” (Art. 3(18) AIA). However, i) the term ‘lifecycle’ is not defined, creating ambiguity about whether it differs from the term ‘lifetime’ used in Art. 12(1) AIA and Rec. (71); ii) the concept of ‘consistent’ performance is unclear, and it is not specified how it should be measured.
First, ‘lifecycle’ and ‘lifetime’ could be understood as synonyms (Marcus, 2020). On the other hand, the term ‘lifetime’ could be understood to refer specifically to the active period of the AI system in operation (Murakami et al., 2010), while ‘lifecycle’ could encompass a broader view of all phases from product design and development to decommissioning (Hamon et al., 2024). In this case, however, it is unclear how accuracy, robustness and cybersecurity should be ensured beyond the operational phase (e.g., during development). Art. 2(8) AIA clarifies that these requirements do not have to be met during the test and development phase of the HRAIS–unless the system is tested under real world conditions. However, the use of the term ’lifecycle’ might be interpreted to suggest that the requirements of Art. 15 AIA should not only be assessed when the system is ready for deployment but also be considered during design process itself.
Second, it is unclear what ‘consistent’ performance means and how it should be measured. In the ML literature, a model’s variability in performance over time is often measured using the variance of a metric such as accuracy or robustness (Kilbertus et al., 2020; Bechavod et al., 2019; Rateike et al., 2022). The variance of a metric over a time interval indicates its deviation from its mean within this interval. For instance, high variance in robustness indicates significant fluctuations in robustness levels between two points in time, whereas low variance indicates similar levels of robustness over time. A low variance could therefore be understood as a consistent performance. [^13] Some also consider consistency as a metric itself, rather than as a property of a (robustness) metric (Wei and Zhang, 2020). In practice, performance can vary due to factors, such as random initializations of weights or input data sampling. These types of variations are unavoidable. Defining level of variance considered ‘consistent’ is challenging as it is dependent on the context. Technical standards and guidelines by the EU Commission should clarify how to measure a consistent performance with respect to accuracy, robustness, and cybersecurity, and provide guidance on determining the required level of consistency.
5.2. Robustness Art. 15(4) AIA
We now turn to challenges specific to Art. 15(4) AIA. Art. 15(4)(i) AIA states that “technical and organisational measures shall be taken” to ensure that AI systems are “as resilient as possible regarding errors, faults or inconsistencies that may occur within the system or the environment”. Art. 15(4)(ii) AIA specifies that robustness can be achieved through technical redundancy solutions, and Art. 15(4)(iii) AIA requires addressing feedback loops in online learning with possibly biased outputs.
Inconsistent Terminology.
The term robustness is used inconsistently throughout the AIA. Art. 15(1) and (4) AIA refer to robustness, whereas the corresponding Rec. (27) and Rec. (75) both mention technical robustness. One could argue that technical robustness is synonymous with robustness. The term ‘technical robustness’ in Rec. (27) may be a remnant of the legislative process that built on the 2019 Ethics Guidelines for Trustworthy AI (AI IHEG, 2019) developed by the AI IHEG, which introduced the principle of ‘technical robustness and safety’. These guidelines are explicitly referenced by Rec. (27). Nevertheless, it remains unclear why Rec. (75) also refers to ‘technical robustness’. It could be that the wording in Rec. (75) is borrowed from Rec. (27). Alternatively, one could argue that robustness in Art. 15(1) and (4) AIA is not limited to technical aspects, but additionally includes some form of non-technical robustness. The latter could refer to organizational measures that must be implemented to ensure robustness (Art. 15(4)(i) AIA). Technical standards and guidelines by the EU Commission should clarify what aspects robustness encompasses.
Required Level of Robustness.
The AIA creates ambiguities regarding the required level of ‘robustness’. Art. 15(1) AIA mandates that AI systems must achieve an “appropriate level” of robustness. Art. 15(4) AIA, however, demands that AI systems shall be “as resilient as possible” to “errors, faults, or inconsistencies”, suggesting a stricter requirement. This discrepancy initially appears ambiguous, as it is unclear whether HRAIS must simply meet an appropriate standard of robustness or strive for the highest possible level. However, the “appropriate” level stated in Art. 15(1) AIA can be understood as a general principle, which is further specified by Art. 15(4) AIA. Therefore, appropriate with respect to robustness is to be understood as ‘as resilient as possible’.
When determining the appropriate level of robustness of a specific HRAIS, the intended purpose of the system and the generally acknowledged state of the art (SOTA) on AI and AI-related technologies must be taken into account (Art. 8(1) AIA). Art. 9(4) AIA acknowledges that one of the objectives of the required risk management is to achieve an “appropriate balance in the implementation of measures to fulfil” requirements. Art. 9(5) AIA further acknowledges the permissibility of a residual risk, meaning that the measures adopted under the risk management system are not expected to eliminate all existing risks, but rather to maintain these residual risks at an ‘acceptable’ level. The risk management system is a continuous iterative process (Art. 9(1) AIA). This means that the appropriate level of robustness of HRAIS must be regularly determined and updated, taking into account its purpose and the SOTA while balancing it with other requirements.
Feedback Loops.
Art. 15(4)(iii) AIA states that AI systems must be explicitly developed in such a way that they “duly address” feedback loops and “eliminate or reduce” the risks associated with them. According to Rec. (67), feedback loops occur when the output of an AI system influences its input in future operations, an under understanding that aligns with the concept as found in the ML literature. Feedback loops are a well-studied problem manifesting in various forms (Pagan et al., 2023), with the most common issues being a distribution shift (Perdomo et al., 2020) or a selection bias (Kilbertus et al., 2020; Lum and Isaac, 2016). Importantly, in this context, the risk of ’biased outputs’ in feedback loops (Art. 15(4)(iii) AIA) is often studied in the literature on fairness in ML rather than in the literature on robustness in ML, which traditionally constitute different research fields and communities (Lee et al., 2021).  [^14] For example, whether there is a trade-off between robustness and fairness, or if both pursue similar goals, remains an active discussion in the ML community (Lee et al., 2021; Xu et al., 2021; Pruksachatkun et al., 2021).
An important aspect of Art. 15(4)(iii) AIA is that it applies specifically to AI systems that learn online. Online learning ML models iteratively learn from a sequence of data and continuously update their parameters over time (Hoi et al., 2021). This adaptiveness is reflected in Art. 3(1) AIA as a factual characteristic of an AI system. The problem with feedback loops in online learning is that newly collected training data can become biased, e.g., due to selection bias, which occurs when the data collected is not representative of the overall population (Zadrozny, 2004; Liu and Ziebart, 2014). This can distort model predictions and reinforce existing biases, ultimately impacting the model’s accuracy and fairness (Kilbertus et al., 2020; Bechavod et al., 2019; Rateike et al., 2022). Offline models, in contrast, are trained on a fixed dataset all at once (Hoi et al., 2021). Offline models can also carry risks when feedback loops are present: The outputs of an ML model can induce a distribution shift through their interaction with the environment (D’Amour et al., 2020; Zhang et al., 2020; Liu et al., 2018). Since an offline ML model is not updated, distribution shifts can influence their performance over time and possibly lead to fairness concerns (Liu et al., 2018). Although Art. 15(4) AIA does not explicitly address feedback loops in offline systems, HRAIS are not exempt from addressing them. Since they can impact the model’s accuracy, feedback loops in offline systems may still need to be addressed to comply with Art. 15(1) AIA.
5.3. Cybersecurity Art. 15(5) AIA
We now turn to legal challenges specific to Art. 15(5) AIA. Art. 15(5)(i) AIA states that AI systems shall be resilient against attempts to “alter their use, outputs, or performance by exploiting system vulnerabilities”. Art. 15(5)(ii) AIA specifies that technical solutions aiming to ensure resilience against such malicious attempts “shall be appropriate to the relevant circumstances and the risks”. Finally, Art. 15(5)(iii) AIA mandates specific measures “to prevent, detect, respond to, and control for attacks” exploiting AI-specific vulnerabilities. This section examines the key aspects of compliance with Art. 15(5) AIA. However, a mentioned above, providers have an additional pathway for demonstrating compliance with its cybersecurity requirements, namely a certification under the CSA (Casarosa, 2022).
Required Level of Cybersecurity.
Art. 15(5)(ii) AIA mandates that technical solutions must be “appropriate to the relevant circumstances and the risks”, but this needs further clarification. The AIA specifically addresses only three kinds of risks: health, safety, and fundamental rights (Rec. (1)). Risks associated with these aspects can be identified and managed through a risk management system that must be put into place as stipulated by Art. 9 AIA. Relevant circumstances are any known and foreseeable circumstances that may have an impact on cybersecurity. [^15] See Art. 13(3)(b)(ii) AIA and Appendix D.
Mandating a cybersecurity level that is ‘appropriate to the relevant circumstances’ acknowledges that complex ML models generally cannot be expected to be fully resistant to all types of adversarial attacks. This has two major reasons: First, it is impossible to anticipate all types of possible attacks. This is acknowledged by Art. 9(5) AIA which states that measures adopted under the risk management system are not expected to remove all existing risks. Second, complete protection against a specific attack cannot be guaranteed, especially as adversaries continuously adapt their strategies to overcome possible defense mechanisms (Xie et al., 2023; Kumar et al., 2023). The appropriateness of a certain performance level must consider the intended purpose of the system and the generally acknowledged SOTA (see Art. 8(1) AIA). The measures to ensure cybersecurity adopted are not expected to eliminate all existing risks, but the overall residual risk must be acceptable (see Art. 9(1) and (4) AIA). Thus, when determining the appropriateness of technical solutions, all applicable requirements of the AIA must be balanced, while also mitigating risks to health, safety, and fundamental rights.
Lastly, while HRAIS are expected to be ‘as resilient as possible’ in terms of robustness, they need only to be ‘resilient’ in terms of cybersecurity. Consequently, the wording of Art. 15 AIA suggests that the robustness requirements are stricter. This may be due to the nature of unintentional causes—such as errors, faults, inconsistencies, or unexpected situations within the system or its operating environment—which are primarily within the provider’s control and justify a higher duty of care. In contrast, attacks by unauthorized third parties are less controllable and therefore justify a (slightly) lower standard for the provider’s duty regarding cybersecurity.
AI-specific Vulnerabilities.
Art. 15(5) AIA differentiates between ’system vulnerabilities’ (Art. 15(5)(i) AIA) and ’AI-specific vulnerabilities’ (Art. 15(5)(iii) AIA). As the term vulnerability is not defined, we provide a working definition. The United States’ Common Vulnerabilities and Exposures (CVE) system defines vulnerability as “[a]n instance of one or more weaknesses […] that can be exploited, causing a negative impact to confidentiality, integrity, or availability” (CVE, 2024). Art. 15(5)(iii) AIA provides a non-exhaustive list of components of an AI system that expose AI-specific vulnerabilities, such as training data, pre-trained components used in training, inputs, or the AI model. However, there might be additional components of the AI system that may also harbor AI-specific vulnerabilities. The question is how to identify them. We suggest performing a hypothetical test. AI models play a central role in an AI system. If a vulnerability would be eliminated by replacing the AI model with a non-AI model, it should be deemed ‘AI-specific’. To define a non-AI model, we return to the definition of an AI system under the AIA. It has been argued that the central characteristic of an AI system is its ability to infer from input to output (Hacker, 2024). This inference ability is typically performed by one or more AI models within an AI system. Therefore, non-AI models are all models lacking inference capability, such as rule-based decision-making systems that rely on predefined rules and logic defined by human experts. [^16] Note that non-AI rule-based systems use human-defined rules, while rule-based ML models infer rules from data (Naik et al., 2023; Weiss and Indurkhya, 1995), qualifying as AI models. In a different context, the AI IHEG ethics guidelines (AI IHEG, 2019) suggest fallback plans where AI systems switch from a statistical (ML) approach to a rule-based or human-in-the-loop approach. Since AI-specific vulnerabilities relate to specific components of an AI system, we suggest viewing them as a subset of system vulnerabilities. To enhance clarity, technical standards should define both terms and mandate a process for identifying them.
Technical Solutions.
Art. 15(5)(iii) AIA provides a non-exhaustive list of attacks and AI-specific vulnerabilities that must be addressed through technical solutions: data poisoning, model poisoning, adversarial examples, model evasion, and confidentiality attacks, which are well-established in the ML literature. These attacks aim to induce model failures (Vassilev et al., 2024): Data poisoning attacks manipulate training data (Schwarzschild et al., 2021), model poisoning attacks manipulate the trained ML model (Zhang et al., 2022), and model evasion attacks manipulate test samples (Biggio et al., 2013). Confidentiality attacks, typically explored in the field of privacy in ML, refer to attempts to extract information about the training data or the model itself (Rigaki and Garcia, 2023).
In addition to these attacks, Art. 15(5)(iii) AIA lists ’model flaws’ as an AI-specific vulnerability. This is a vague legal term and lacks an established counterpart in the ML literature. In software contexts, the word flaw often refers to so-called bugs, which are typically the result of human errors in the coding process (Kumar and Anderson, 2023; Nissenbaum, 1996). However, the term model flaw follows the list of attacks outlined above, which are instead designed to exploit the default properties of a properly functioning ML model, and are not directly the results of errors in the coding process. Thus, it is unclear what model flaw refers to in this context, and whether technical solutions are only expected to address traditional bugs or coding errors, or whether they should address other ways of exploiting AI-specific vulnerabilities. Given that the term is situated within the cybersecurity requirements for AI system outlined in Art. 15(5) AIA, we argue that the term model flaws should be interpreted as flaws that enable the exploitation of AI-specific vulnerabilities. Technical standards and guidelines by the EU Commission should define model flaws more clearly and provide guidelines for technical solutions to address these model flaws. This should take into account the arms race between attacker and defender in the realm of adversarial robustness, where both parties are continuously adapting their strategies to outmaneuver the other (Chen et al., 2017b). This makes it infeasible to anticipate and counter all potential attacks that target AI-specific vulnerabilities.
Organizational Measures.
Numerous EU regulations related to cybersecurity (see e.g., Art. 32 General Data Protection Regulation [^17] EU Regulation 2016/679, OJ L 119, 4.5.2016., Art. 21 NIS 2 Directive [^18] EU Directive 2022/2555, OJ L 333/80.) explicitly mandate both technical and organizational measures to ensure cybersecurity. In the AIA, organizational measures are only mandated for the robustness of HRAIS in Art. 15(4) AIA, but not for cybersecurity (Art. 15(5) AIA). Rather, Art. 15(5) AIA AIA only focuses on technical solutions for providers of HRAIS. The omission of organizational measures in Art. 15(5) AIA has been criticized in the literature accompanying the legislative process of the AIA (Biasin et al., 2023). It is unclear whether providers are still implicitly required to implement organizational measures (in accordance with other EU regulations), as these measures might be inherently included in the concept of cybersecurity, or if they are not mandatory. However, this ambiguity for providers of HRAIS (Art. 15 AIA) is mitigated by the fact that deployers of HRAIS are required to implement both organizational and technical measures to ensure the proper use of the system in accordance with the instructions for use (Art. 26(1) AIA). These instructions include the cybersecurity measures put in place.
6. Requirements for General-Purpose AI Models With Systemic Risk
In the previous section, we examined HRAIS requirements. To further elucidate them, we study GPAIMs with systemic risk, highlighting similarities and differences. The AIA establishes legal requirements for GPAIMs, such as multimodal large language models (OpenAI, 2023; Gemini Team et al., 2023), which can perform tasks beyond their original training objective (Gutierrez et al., 2023). GPAIM can be standalone or embedded in an HRAIS, with the latter requiring compliance with both GPAIM and HRAIS requirements. The AIA distinguishes between GPAIM with systemic risks and those without. Art. 3(65) AIA defines ‘systemic risk’ as the risk that is specific to the high-impact capabilities of GPAIMs that have a “significant impact” on the market, public health, safety, security, fundamental rights, or society. [^19] A systemic risk is presumed when the cumulative computation during training exceeds $10^{25}$ Floating-Point Operations Per Second (FLOPS). GPAIMs with fewer FLOPS may still be classified as posing a systemic risk under Art. 51(1) AIA. There is an ongoing debate over this threshold and if a model’s complexity truly reflects its risk level (Hacker, 2024; Novelli et al., 2024; Pehlivan, 2024; Kutterer, 2023). Thresholds and criteria can be modified by the EU Commission, see Appendix B GPAIMs without systemic risks are exempt from robustness and cybersecurity obligations (Art. 53 AIA ff.).
Cybersecurity Requirements.
Art. 55(1)(d) AIA mandates “an adequate level of cybersecurity protection” for GPAIMs with systemic risk. Rec. (115) further details this cybersecurity requirement. It mandates cybersecurity protection against “malicious use or attacks” and lists specific adversarial threats, such as “accidental model leakage, unauthorised releases, circumvention of safety measures”, “cyberattacks”, or “model theft”. Notably, several of these threats have direct counterparts in the ML literature on adversarial robustness and privacy for large generative models, such as the circumvention of safety measures (jailbreaking) or model theft (Yao et al., 2024; Li et al., 2023; Wang et al., 2023). Although Art. 55(1)(d) AIA does not define the term ‘cyberattacks’, we infer that it includes the attacks exploiting AI-specific vulnerabilities mentioned in Art. 15 AIA (see Section 5.3). These attacks are studied in the field of adversarial robustness and can also affect GPAIMs (Qiang et al., 2024; Das et al., 2024; Yan et al., 2024; Schwinn et al., 2024; Vitorino et al., 2024)—even though specific ML techniques may be necessary to address GPAIM-specific challenges. This relation underscores that the concepts and problems explored under adversarial robustness are reflected in the term ’cybersecurity’ as used in Art. 55(1)(d) AIA. To ensure the cybersecurity of a GPAIMs with systemic risk, providers must conduct and document internal and/or external adversarial testing of the model, such as red teaming. [^20] In this context, red teaming refers to stress testing AI models by simulating adversarial attacks (Feffer et al., 2024), such as linguistic or semantic attacks against LLMs (Shi et al., 2024), whereas traditional cybersecurity red teaming focuses on assessing entire systems or networks (Teichmann and Boticiu, 2023).
While the AIA mandates robustness requirements for HRAIS, we observe that it does not impose an explicit equivalent legal requirement for GPAIMs, regardless of whether they present a systemic risk or not. Specifically, neither Art. 55 AIA nor Rec. (155) address unintentional causes for deviations from consistent performance. In Section 5.2, we stated that non-adversarial robustness is reflected in the term robustness in Art. 15 AIA. Consequently, GPAIMs, which are not required to fulfill any robustness requirement, are not mandated to be resilient against performance issues, such as data distribution shifts or noisy data. The AIA itself does not provide an explanation for the omission of a robustness requirement. It may stem from the complexity of political negotiations regarding the AIA, particularly regarding GPAIMs, which were not addressed in the initial draft of the regulation but gathered widespread media attention during the legislative procedure. However, evidence from ML research suggests that non-adversarial robustness is also relevant for GPAIMs (Yuan et al., 2023; Chen et al., 2022).
Required Level of Cybersecurity.
Art. 55(1)(d) AIA mandates an ‘adequate’ level of cybersecurity protection for GPAIMs with systemic risk. This requirement contrasts with the ‘appropriate’ level of cybersecurity mandated for HRAIS under Art. 15(1) AIA. The use of these two different terms raises questions about whether both HRAIS and GPAIMs should achieve the same level of cybersecurity or to what extent their required levels might differ. On the one hand, ‘adequate’ and ‘appropriate’ could imply different levels of cybersecurity. The Cambridge Dictionary defines the term ‘adequate’ as “enough or satisfactory for a particular purpose” (Cambridge University Press, 2024a) and ‘appropriate’ as “suitable or right for a particular situation or occasion” (Cambridge University Press, 2024b). Accordingly, something is ‘adequate’ if it exceeds a minimum threshold that is good enough, while something is ‘appropriate’ if it meets a specific (right) level above that minimum. GPAI models can perform a wide variety of tasks in different contexts and thus be prone to a variety of different intentional causes of harm, making it difficult to identify and mitigate their specific cybersecurity risks. For this reason, it may be reasonable to only mandate an ‘adequate’, i.e., minimum level of cybersecurity. HRAIS, independently of whether they contain an GPAIM as a component, can be thought of as operating in a more specific contexts, potentially allowing an easier and more precise assessment of cybersecurity risks and thus a more stringent appropriate level of cybersecurity protection. On the other hand, ‘adequate’ and ‘appropriate’ could refer to the same level of cybersecurity. Rec. (115) states that “adequate technical and established solutions” must be “appropriate to the relevant circumstances and the risks”. The simultaneous use of both terms in a single sentence, intended to guide the interpretation of Art. 55(1)(d) AIA, suggests that they might be intended as synonymous. This is corroborated by the observation that many official language versions of the AIA use a single term for both “adequate” and “appropriate” in Art. 15(1) AIA and Art. 55(1)(d) AIA. [^21] Such as FR “approprié”, ES “adecuado”, GER “angemessen”, IT “adeguato”. To resolve this ambiguity, technical standards should clarify the required level of cybersecurity for GPAIMs with systemic risk.
7. Summary and Outlook
We identified several legal challenges and potential limitations in implementing robustness and cybersecurity requirements for HRAIS under Art. 15(4) and (5) AIA. We also examined GPAIMs with systemic risk, which face cybersecurity but not robustness requirements, and identified additional legal challenges. We proposed a simple explanatory model that maps non-adversarial robustness in the ML literature to the term ‘robustness’ used in Art. 15(1) and (4) AIA, and that maps adversarial robustness to the term ‘cybersecurity’ used in Art.   AIA15(1) and (5) AIA. However, both ‘robustness’ and ‘cybersecurity’ can refer also to other concepts both within the domain of ML and beyond. Comparing the provisions for HRAIS to those for GPAIMs with systemic risk, we argued that adversarial robustness maps to the term ‘cybersecurity’ used in Art. 55(1)(d) AIA. However, we were not able to find an explicit equivalent legal requirement for non-adversarial robustness in the provisions regulating GPAIMs with systemic risk models.
Our analysis highlights the need for clearer specifications of these provisions through harmonized standards, guidelines by the EU Commission or the benchmark and measurement methodologies foreseen for robustness and cybersecurity in Art. 15(2) AIA. These would help define technical requirements and establish evaluation criteria for AI systems. Specifically, we suggest that technical standards and guidelines by the EU Commission should focus on: i) identify the technical requirements associated with vague legal terms; ii) defining the required level of ‘robustness’ and ‘cybersecurity’ and other concepts such as ‘consistency’; iii) defining the requirements for evaluating and assessing AI systems and its components; and iv) pay attention to some aspects that are not explicitly regulated, such as feedback loops in offline systems in Art. 15(4) AIA or organizational measures to ensure ‘cybersecurity’ in Art. 15(5) AIA.
However, while standards and guidelines can ensure compatibility and practical integration of regulatory frameworks, they can struggle to keep pace with rapid technological advancements. This can lead to outdated versions that do not fully address emerging technologies or novel applications. Our analysis is not without limitations. Due to the novelty of the AIA, we lack empirical data to support claims about the challenges in implementing its requirements. While we focus on identifying these challenges, proposing specific definitions, processes, metrics, or thresholds is left for future work. Future research should focus on non-adversarial robustness for GPAIMs with systemic risk, and explore legal intersections with frameworks like the Medical Device Regulation (Biasin et al., 2023; Nolte and Schreitmüller, 2024). Additionally, the focus on models in ML research versus entire AI systems in the AIA underscores the need for interdisciplinary work. Within the ML domain, future work should explore the impact of accuracy metrics on robustness, potential ‘robustness hacking’, and methods to measure and ensuring long-term performance consistency in the presence of feedback loops.
Appendix
Appendix A Legal Terminology
The AIA is formally structured into recitals (Rec.), articles (Art.), and annexes. Recitals are legally non-binding and outline the rationale behind the articles, articles delineate specific binding obligations, and the annexes provide additional details and specifications to support the articles (Klimas and Vaiciukaite, 2008).
Appendix B Notions of High-Risk AI Systems and General Purpose Models provided in AIA
AI System.
The AIA defines an AI system as “a machine-based [^22] Thereby, machine-based “refers to the fact that AI systems run on machines” (Rec. (12)). system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments” (Art. 3(1) AIA). Thereby, Rec. (12), suggests that notion of ‘AI system’ “should be clearly defined and should be closely aligned with the work of international organisations working on AI to ensure legal certainty, facilitate international convergence and wide acceptance, while providing the flexibility to accommodate the rapid technological developments in this field”. Importantly, “the definition should be based on key characteristics of AI systems that distinguish it from simpler traditional software systems or programming approaches and should not cover systems that are based on the rules defined solely by natural persons to automatically execute operations” (Rec. (12)). Thereby a “key characteristic of AI systems is their capability to infer”, which refers to “the process of obtaining the outputs, such as predictions, content, recommendations, or decisions, which can influence physical and virtual environments, and to a capability of AI systems to derive models or algorithms, or both, from inputs or data” (Rec. (12)). Examples for techniques that enable inference “include machine learning approaches that learn from data how to achieve certain objectives, and logic- and knowledge-based approaches that infer from encoded knowledge or symbolic representation of the task to be solved” (Rec. (12)).
High-risk AI Systems (HRAIS).
The AIA classifies AI systems into different risk groups. Risk refers thereby to “the combination of the probability of an occurrence of harm and the severity of that harm” (Art. 3(2) AIA). An AI system is considered high-risk, if it is “intended to be used as a safety component of a product, or the AI system is itself a product” and “is required to undergo a third-party conformity assessment, with a view to the placing on the market or the putting into service of that product” covered by the Union harmonisation legislation listed in Annex I (Art. 6(1) AIA). Annex I AIA provides a list of 20 EU harmonisation legislation. For example, the Machinery Directive (Directive 2006/42/EC and amending Directive 95/16/EC), the directive on the safety of toys (Directive 2009/48/EC), or the directive concerning agricultural and forestry vehicles (Regulation (EU) No 167/2013). This means that for example, an AI system that is used in a product or as a product itself considered as a toy falling under Directive 2009/48/EC, would be considered as a high-risk AI system.
An AI systems is also considered high-risk, if it operates in types or use-cases enlisted in Annex III (Art. 6(2) AIA) and poses a significant risk of harm to the health, safety or fundamental rights of natural person following (Art. 6(3) AIA). Annex III enlists eight types or use-cases of AI systems: biometric applications (e.g., remote biometric identification systems excluding biometric verification, biometric categorisation, and emotion recognition); critical infrastructure (e.g., “critical digital infrastructure, road traffic, or in the supply of water, gas, heating or electricity”); education and vocational training; employment, workers management and access to self-employment; essential private and public services; law enforcement; migration, asylum and border control management; and administration of justice and democratic processes. Thereby, the Commission shall “provide […] a comprehensive list of practical examples of use cases of AI systems that are high-risk and not high-risk” (Art. 6(5) AIA). Rec. (52) suggests that “it is appropriate to classify them [i.e., AI systems] as high-risk if, in light of their intended purpose, they pose a high risk of harm to the health and safety or the fundamental rights of persons, taking into account both the severity of the possible harm and its probability of occurrence and they are used in a number of specifically pre-defined areas”. Thereby the methodology and citeria for the identification of high-risk should be able to be adopted in order to account for “the rapid pace of technological development, as well as the potential changes in the use of AI systems” (Rec. (52)). For example, an AI system used for remote biometric identification is considered high-risk, unless it does not pose a significant risk. Remote identification refers to comparing biometric data of that individual to stored biometric data of individuals in a reference database (Rec. (15)), where a remote biometric identification system should be understood as a “AI system intended for the identification of natural persons without their active involvement, typically at a distance” (Rec. (17)).
General-purpose AI Models (GPAIM).
A general-purpose AI model (GPAIM) refers to an “AI model […] that displays significant generality and is capable of competently performing a wide range of distinct tasks regardless of the way the model is placed on the market and that can be integrated into a variety of downstream systems or applications” (Art. 3(63) AIA). The AIA, however, explicitly excludes from its regulation AI models that may fall under this definition but are “used for research, development or prototyping activities before they are placed on the market” (Art. 3(63) AIA). Rec. (97) suggest that the term general-purpose AI model “should be clearly defined and set apart from the notion of AI systems to enable legal certainty” taking into account the “key functional characteristics of a general-purpose AI model, in particular the generality and the capability to competently perform a wide range of distinct tasks”. Thereby, GPAIMs “may be placed on the market in various ways, including through libraries, application programming interfaces (APIs), as direct download, or as physical copy”, and “may be further modified or fine-tuned into new models” (Rec. (97)). Examples for GPAIMs, are large generative AI models that “allow for flexible generation of content, such as in the form of text, audio, images or video, that can readily accommodate a wide range of distinctive tasks” (Rec. (99)). This is the case for many multimodal large language models, such as GPT-4 Omni (GPT-4o)  [^23] https://openai.com/gpt-4o-contributions/, or Gemini  [^24] https://gemini.google.com.
GPAIM with Systemic Risk.
The systemic risk of a GPAIM is understood as “a risk that is specific to the high-impact capabilities of general-purpose AI models, having a significant impact on the Union market due to their reach, or due to actual or reasonably foreseeable negative effects on public health, safety, public security, fundamental rights, or the society as a whole, that can be propagated at scale across the value chain” (Art. 3(65) AIA). Rec. (110) provides as a non-exhaustive list of examples for systemic risks: “any actual or reasonably foreseeable negative effects in relation to major accidents, disruptions of critical sectors and serious consequences to public health and safety; any actual or reasonably foreseeable negative effects on democratic processes, public and economic security; the dissemination of illegal, false, or discriminatory content”. Thereby “[s]ystemic risks should be understood to increase with model capabilities and model reach, can arise along the entire lifecycle of the model, and are influenced by conditions of misuse, model reliability, model fairness and model security, the level of autonomy of the model, its access to tools, novel or combined modalities, release and distribution strategies, the potential to remove guardrails and other factors” (Rec. (110)).
According to Art. 51(1)(a) AIA, a GPAIM is considered posing a systemic risk, if “it has high impact capabilities evaluated on the basis of appropriate technical tools and methodologies, including indicators and benchmarks”. High-impact capabilities are understood as “capabilities that match or exceed the capabilities recorded in the most advanced general-purpose AI models” (Art. 3(64) AIA). Thereby high impact capabilities of a GPAIM are to be assumed, “when the cumulative amount of computation used for its training measured in floating point operations is greater than $10^{25}$” (Art. 51(2) AIA). Rec. (111) states that “according to the state of the art at the time of entry into force of this Regulation, the cumulative amount of computation used for the training of the general-purpose AI model measured in floating point operations is one of the relevant approximations for model capabilities”. According to Art. 51(1)(b) AIA a GPAIM is also considered to pose a systemic risk if—“based on a decision of the Commission”—it is considered having capabilities or an impact equivalent to those set out in Art. 51(1)(a) AIA according to the criteria set out in Annex XIII. The seven criteria enlisted in Annex XIII include the number of model parameters, the quality or size of the data set, the amount of computation used for training the model, the input and output modalities of the model, the benchmarks and evaluations of capabilities of the model, whether it has a high impact on the internal market due to its reach, and the number of registered end-users.
Importantly, according to Art. 51(3) AIA, the EU Commission shall be able to amend the thresholds in Art. 51(1) and (2) AIA and add “benchmarks and indicators in light of evolving technological developments, such as algorithmic improvements or increased hardware efficiency, when necessary, for these thresholds to reflect the state of the art” (see also Rec. (111)).
Appendix C Lifecycle in Art. 15(1) AIA
The exact time frame during which consistent performance must be ensured following Art. 15(1) AIA is unclear. Particularly, the term ‘lifecycle’ is not defined, leaving open whether it differs from the term ‘lifetime’ used in Art. 12(1) AIA and Rec. (71). It is crucial to clarify the exact timeframe during which consistent performance must be maintained. While ‘lifecycle’ and ‘lifetime’ could initially be interpreted as synonyms (Marcus, 2020), ‘lifetime’ might refer specifically to the active operational period of the AI system (Murakami et al., 2010), whereas ‘lifecycle’ could encompass a broader view of all phases from product design and development to decommissioning (Hamon et al., 2024). If this broader interpretation of ‘lifecycle’ is intended, it raises questions about how accuracy, robustness, and cybersecurity should be ensured beyond the operational phase (e.g., during development), and why this would be necessary when there are no immediate risks to health, safety, and fundamental rights. One explanation for using the term ’lifecycle’ would be that the EU legislator intended to emphasize that the requirements of Art. 15 AIA should not only be assessed when the system is ready for deployment but also during the design process. Accordingly, technical standards should define both terms.
Appendix D Relevant Circumstance Art. 15(5)(ii) AIA
The term ’relevant circumstance’ is not defined in the AIA and therefore requires interpretation. On the one hand, one could argue that the term only refers to circumstances that are “important” for a “particular purpose” or context (Cambridge University Press, 2024c). On the other hand, the meaning of the term can also result from a comparison with other provisions of the AIA such as Art. 13(3)(b)(ii) AIA, which suggests a different understanding. The provision demands that the instructions for the use of AI systems shall contain “any known and foreseeable circumstances” that may have an impact on cybersecurity. This speaks for a broader understanding of relevance, which only excludes unknown and unforeseeable circumstances. Given this ambiguity, standards should elaborate on how to determine relevant circumstances.
Appendix E Robustness-Accuracy Trade-Off
The ML literature has found that robustness and accuracy of ML models can in some scenarios be empirically and theoretically mutually inhibiting (Zhang et al., 2019; Li and Li, 2024; Yang et al., 2020; Raghunathan et al., 2020; Rade and Moosavi-Dezfooli, 2022), and this relationship remains an active area of research (Pang et al., 2022). As stated in Section 5.1, an AI system may also incorporate other technical components (beyond models) that must be robust and may impact the overall system’s accuracy. The AIA acknowledges these trade-offs but does not offer specific guidelines on how to achieve this balance. It requires providers of HRAIS to ensure an appropriate level of both accuracy and robustness, and to find “an appropriate balance in implementing the measures to fulfil” the AIA requirements (Art. 9(4) AIA). Additionally, the technical documentation must include “decisions about any possible trade-off made regarding the technical solutions adopted to comply with the requirements” of the AIA (Annex IV Nr. 2 lit. b AIA). Standards may offer guidance on the processes and metrics to use but will not prescribe a specific balance, such as a 40-60 ratio. As a result, providers of HRAIS will ultimately need to navigate these complex trade-offs themselves, adjusting individual model parameters to find an appropriate balance and justify and document their decisions.